find foou conscutive numbers that add to 44



Answer by checkley77(12844)

I'll assume that foou is four.
x+x+1+x+2+x+3=44
4x+6=44
4x=44-6
4x=38
x=38/4
x=9.5 first number
9.5+1=10.5 second number
9.5+2=11.5 third number
9.5+3=12.5 fourth number.
Proof:
9.5+10.5+11.5+12.5=44
44=44